# RCA: Region Conditioned Adaptation for Visual Abductive Reasoning

### v4

@@ Abstract.
Refer to caption (a)
Refer to caption (b)
@@
Visual Abductive Reasoning, CLIP Adapter, Region Prompt Tuning
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 28 - November 1, 2024; Melbourne, Australia.isbn: 978-1-4503-XXXX-X/18/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
1. Introduction
@@
Recently, a novel multimodal visual reasoning problem known as Visual Abductive Reasoning (VAR) has been introduced (Hessel et al., 2022), highlighting the significance of integrating both visual and textual modalities to infer logical conclusions from observed image data. Visual abductive reasoning refers to making inferences based on visual information (usually an incomplete set of observations) to arrive at the most plausible (often simplest) explanation or hypothesis for a given observation. For example, in VAR, as shown in LABEL:Fig:task_define, the model is expected to make the inference “the woman recently ordered the drink” from the given regional visual hints, which show only the “glass bottle” and surrounding contexts ( “restaurant scene” and “waitress”). Visual abductive reasoning is challenging because it requires a deep understanding of the observed image and the domain (or the context) of the scene depicted in the image. Furthermore, VAR demands the ability to generate hypotheses consistent with the observed visual data and the domain rules. It involves not only recognizing patterns in the images but also applying domain knowledge, learning to reason about unseen/unknown aspects of the context, and learning the causal relationship between inferences and observations.
@@ 2. Related Works
Foundation Models. Scaling up models’ complexities and training data improves the attention-based (Galassi et al., 2020; Otter et al., 2020; Liu et al., 2023; Hao et al., 2022a) foundation models’ (Brown et al., 2020; Devlin et al., 2019; Radford et al., 2021; Jia et al., 2021; Yu et al., 2022; Yuan et al., 2021) perception capacity, making it proficient in many tasks including zero or few-shot learning. Specifically, Large Language Models (LLM), such as BERT (Devlin et al., 2019), and GPT (Brown et al., 2020) are trained on large-scale datasets and they generalize to many downstream NLP tasks. Following this trend, several vision-language foundational models are also developed e.g. CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021) and BLIP (Li et al., 2022). The main idea behind the vision foundation models is to learn transferable visual representation with noisy text supervision through a two-tower structure. We follow the current baseline of visual abductive reasoning and adopt the CLIP model as the backbone for visual inference.
@@
Parameter-Efficient Fine-Tuning (PEFT). Transferring foundational models to downstream tasks promotes the development of PEFTs (Sabry and Belz, 2023). Representative PEFTs include Prompt, Adapter, and LoRA tuning. Specifically, prompt tuning (Brown et al., 2020; Lester et al., 2021; Zhong et al., 2021; Tu et al., 2022; Liu et al., 2022; Ma et al., 2024; Wang et al., 2024) enhances the distinctiveness of inputs by prepending additional tokens, which may be either trainable or fixed. The vision-language prompt tuning can further be divided into textual (Zhou et al., 2022b, a; Feng et al., 2022; Du et al., 2022) or visual prompt (Jia et al., 2022; Bahng et al., 2022) tuning, depending on the placement of prompt tokens in visual or textual encoders. Certain special visual prompts, such as the Merlot (Zellers et al., 2021), CPT (Yao et al., 2024), and CiP (Shtedritski et al., 2023), guide the model to focus on specified areas by overlaying these regions with translucent colours or red circles. In adapter tuning, trainable Multi-Layer Perceptron (mini MLP) (Yang et al., 2023; Sung et al., 2022; Houlsby et al., 2019) or Tiny Attention modules (Zhao et al., 2022b) are usually inserted into the foundational model, with only the new additions being fine-tuned. LoRA (Hu et al., 2022) update parameters using low-rank projections. Our RCA is a hybrid “prompt+adapter” tuning to equip the vision foundational models with local reasoning ability, an approach that has not been studied before.
@@ 3.3. Regional Prompt Generator
In visual abductive reasoning, it is important to collect all relevant visual cues from the image and the region $\boldsymbol{r}$. Therefore, we use pre-specified observation $\boldsymbol{r}$ as a “prompt” to guide the visual reasoning process directly. Our Regional Prompt Generator (RPG) creates three detailed prompts focusing on specific regions. These prompts harness local features (i.e., region), surrounding context, and existing visual prompts, as shown in Figure LABEL:fig:prompt1-LABEL:fig:prompt3. All three types of prompts go through the same process, with the only difference in being whether colors or circles (Shtedritski et al., 2023) are drawn on the input images. To explain how it works, we’ll use the “Region+Context” (R-CTX) as an example.
@@ 3.4. Adapter+ Tuning
(6) $\boldsymbol{z^{\prime\prime}}_{l}={\color[rgb]{.75,0,.25}\text{Adapter\_A}}\left(\boldsymbol{z^{\prime}}_{l}\right)+\boldsymbol{z}_{l-1},$
(7) $\boldsymbol{z}_{l}={\text{MLP}}\left(\boldsymbol{z^{\prime\prime}}_{l}\right)+{\color[rgb]{.75,0,.25}\text{Adapter\_M}}\left(\boldsymbol{z^{\prime\prime}}_{l}\right)+\boldsymbol{z^{\prime\prime}}_{l},$
@@
(8) $\boldsymbol{z^{\prime}}_{l}={\text{Softmax}}\left(\frac{\boldsymbol{Q}\boldsymbol{K}^{T}+{\color[rgb]{.75,0,.25}\boldsymbol{\widehat{Q}}\boldsymbol{\widehat{K}}^{T}}}{{\sqrt{D}}}\right)\boldsymbol{V},$
(9) ${\color[rgb]{.75,0,.25}\boldsymbol{\widehat{Q}},\boldsymbol{\widehat{K}}}~{}=\boldsymbol{z}_{l-1}\times{\color[rgb]{.75,0,.25}\boldsymbol{\widehat{W}}_{q},\boldsymbol{\widehat{W}}_{k}},~{}~{}~{}{\color[rgb]{.75,0,.25}\widehat{\boldsymbol{W}}_{q,k}}\in\mathbb{R}^{D\times d}$

### v5

@@ Abstract.
1. Introduction
Visual Abductive Reasoning, CLIP Adapter, Region Prompt Tuning
copyright: acmlicensedjournalyear: 2024copyright: licensedothergovconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiaisbn: 979-8-4007-0686-8/24/10doi: 10.1145/3664647.3681244ccs: Computing methodologies Reasoning about belief and knowledgeccs: Computing methodologies Ranking
@@ 1. Introduction
Recently, a novel multimodal visual reasoning problem known as Visual Abductive Reasoning (VAR) has been introduced (Hessel et al., 2022), highlighting the significance of integrating both visual and textual modalities to infer logical conclusions from observed image data. Visual abductive reasoning refers to making inferences based on visual information (usually an incomplete set of observations) to arrive at the most plausible (often simplest) explanation or hypothesis for a given observation. For example, in VAR, as shown in Figure LABEL:Fig:task_define, the model is expected to make the inference “the woman recently ordered the drink” from the given regional visual hints, which show only the “glass bottle” and surrounding contexts ( “restaurant scene” and “waitress”). Visual abductive reasoning is challenging because it requires a deep understanding of the observed image and the domain (or the context) of the scene depicted in the image. Furthermore, VAR demands the ability to generate hypotheses consistent with the observed visual data and the domain rules. It involves not only recognizing patterns in the images but also applying domain knowledge, learning to reason about unseen/unknown aspects of the context, and learning the causal relationship between inferences and observations.
@@ 2. Related Works
Foundation Models. Scaling up models’ complexities and training data improves the attention-based (Galassi et al., 2020; Otter et al., 2020; Liu et al., 2023; Hao et al., 2022a) foundation models’ (Brown et al., 2020; Devlin et al., 2019; Radford et al., 2021; Jia et al., 2021; Yu et al., 2022; Yuan et al., 2021; Cheng et al., 2024) perception capacity, making it proficient in many tasks including zero or few-shot learning. Specifically, Large Language Models (LLM), such as BERT (Devlin et al., 2019), and GPT (Brown et al., 2020) are trained on large-scale datasets and they generalize to many downstream NLP tasks. Following this trend, several vision-language foundational models are also developed e.g. CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021) and BLIP (Li et al., 2022). The main idea behind the vision foundation models is to learn transferable visual representation with noisy text supervision through a two-tower structure. We follow the current baseline of visual abductive reasoning and adopt the CLIP model as the backbone for visual inference.
@@
Parameter-Efficient Fine-Tuning (PEFT). Transferring foundational models to downstream tasks promotes the development of PEFTs (Sabry and Belz, 2023). Representative PEFTs include Prompt, Adapter, and LoRA tuning. Specifically, prompt tuning (Brown et al., 2020; Lester et al., 2021; Zhong et al., 2021; Tu et al., 2022; Liu et al., 2022; Ma et al., 2024; Wang et al., 2024) enhances the distinctiveness of inputs by prepending additional tokens, which may be either trainable or fixed. The vision-language prompt tuning can further be divided into textual (Zhou et al., 2022b, a; Feng et al., 2022; Du et al., 2022; Fang et al., 2024) or visual prompt (Jia et al., 2022; Bahng et al., 2022; Oh et al., 2024) tuning, depending on the placement of prompt tokens in visual or textual encoders. Certain special visual prompts, such as the Merlot (Zellers et al., 2021), CPT (Yao et al., 2024), and CiP (Shtedritski et al., 2023), guide the model to focus on specified areas by overlaying these regions with translucent colours or red circles. In adapter tuning, trainable Multi-Layer Perceptron (mini MLP) (Yang et al., 2023; Sung et al., 2022; Houlsby et al., 2019) or Tiny Attention modules (Zhao et al., 2022b) are usually inserted into the foundational model, with only the new additions being fine-tuned. LoRA (Hu et al., 2022) update parameters using low-rank projections. Our RCA is a hybrid “prompt+adapter” tuning to equip the vision foundational models with local reasoning ability, an approach that has not been studied before.
@@ 3.3. Regional Prompt Generator
In visual abductive reasoning, it is important to collect all relevant visual cues from the image and the region $\boldsymbol{r}$. Therefore, we use pre-specified observation $\boldsymbol{r}$ as a “prompt” to guide the visual reasoning process. Our Regional Prompt Generator (RPG) creates three detailed prompts focusing on specific regions. These prompts harness local features ($\boldsymbol{r}$), surrounding context, and existing visual prompts, as shown in Figure LABEL:fig:prompt1-LABEL:fig:prompt3. All three types of prompts go through the same process, with the only difference in being whether colors or circles (Shtedritski et al., 2023) are drawn on the input images. To explain how it works, we’ll use the “Region+Context” (R-CTX) as an example.
@@ 3.4. Adapter+ Tuning
(6) $\boldsymbol{z^{\prime\prime}}_{l}={\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\text{Adapter\_A}}\left(\boldsymbol{z^{\prime}}_{l}\right)+\boldsymbol{z}_{l-1},$
(7) $\boldsymbol{z}_{l}={\text{MLP}}\left(\boldsymbol{z^{\prime\prime}}_{l}\right)+{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\text{Adapter\_M}}\left(\boldsymbol{z^{\prime\prime}}_{l}\right)+\boldsymbol{z^{\prime\prime}}_{l},$
@@
(8) $\boldsymbol{z^{\prime}}_{l}={\text{Softmax}}\left(\frac{\boldsymbol{Q}\boldsymbol{K}^{T}+{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\boldsymbol{\widehat{Q}}\boldsymbol{\widehat{K}}^{T}}}{{\sqrt{D}}}\right)\boldsymbol{V},$
(9) ${\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\boldsymbol{\widehat{Q}},\boldsymbol{\widehat{K}}}~{}=\boldsymbol{z}_{l-1}\times{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\boldsymbol{\widehat{W}}_{q},\boldsymbol{\widehat{W}}_{k}},~{}~{}~{}{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\widehat{\boldsymbol{W}}_{q,k}}\in\mathbb{R}^{D\times d}$
@@ 4.3. Comparision with the State-of-the-Art
[…]
×$\times 0$ (Zero-Shot, R-CTX)
00.00 159.31 141.37 7.58 36.75 / 11.88 6.78
[…]
×$\times 0$ (Full Fine-Tuning)
149.77 15.95 17.93 33.55 85.84 / 40.41 26.03